# Online Learning of Interaction Dynamics with Dual Model Predictive Control for Multi-Agent Systems Using Gaussian Processes

### v1

@@ III-C Gaussian Process Prediction Model
We exploit the joint probability of the GP $d^{j}$ to account for the uncertainty in our safety constraints $\mathbb{X}^{j}$ by formulating chance constraints that should be satisfied with a probability $p_{x}\in\left(0,1\right]$. To this end, we optimize the control sequence of Agent 0 $U_{k}=\bigl{(}u^{0}_{0\mid k},\ldots,u^{0}_{N-1\mid k}\bigr{)}$ in the following SOCP:
@@
Note that the posterior GP $d^{j}$ is a function of the control sequence. As this SOCP (10) is generally intractable, we subsequently apply tractable approximations for consecutive evaluations of the GP and the chance constraints, at the expense of formal safety guarantees. In order to propagate the covariance of the predicted state $\Sigma^{x^{j}}$, we make the following assumptions. Firstly, we assume that consecutive GP evaluations are independent. Furthermore, the predicted state $x^{j}$ and the posterior GP $d^{j}$ are assumed jointly Gaussian:
$\begin{bmatrix}x^{j}\\ d^{j}\end{bmatrix}\sim\mathcal{N}\left(\mu^{j},\Sigma^{j}\right)\\ =\mathcal{N}\left(\begin{bmatrix}\mu^{x^{j}}\\ \mu^{d^{j}}\end{bmatrix},\begin{bmatrix}\Sigma^{x^{j}}&\Sigma^{x^{j}d^{j}}\\ \Sigma^{d^{j}x^{j}}&\Sigma^{d^{j}}\end{bmatrix}\right),$ (11)
such that the covariance can be approximated and propagated over the prediction horizon. We approximate $\mu^{d^{j}},\Sigma^{x^{j}{d^{j}}},\Sigma^{d^{j}}$ through a first-order Taylor approximation of the GP evaluated at its posterior mean [14], which provides a good trade-off between accuracy and computational complexity [8]:
@@
for $i=0,1,\dots,N-1$ where the initial prediction equals the current state $\mu^{x^{j}}_{0\mid k}=x^{j}_{k}$ with ${\Sigma^{x^{j}}_{0\mid k}=\mathbf{0}}$. We approximate the full GP by a Sparse Pseudo-Input GP (SPGP) [15], which exploits the Gaussian structure to reduce the computational complexity of the GP by preconditioning the GP on heuristic inducing points. For details on SPGPs, we refer the reader to [15]. We select the inducing points by uniformly sampling from previously predicted trajectory from the MPC, since the test points can be expected to be close to this trajectory [8].
@@ III-D Learning-based Dual Model Predictive Control
In this section, we propose our dual MPC that uses the GP prediction model, presented in Section III-C, to jointly condition the interactions of the other agents with the control inputs of the Agent 00. The dual control effect was first formalized in [3], and according to [4] is defined as follows:
@@
where the chance constraints from (10d) are approximated by a general tightened constraint set $\tilde{\mathbb{X}}^{j}\bigl{(}\mu^{x^{j}},\Sigma^{x^{j}}\bigr{)}\subseteq\mathbb{X}^{j}\bigl{(}\mu^{x^{j}}\bigr{)}$, based on the covariance of the GP [8]. The specific form of constraint tightening may depend on the nature of the problem. An example hereof is provided in Section IV. The dual MPC, detailed in Algorithm 1, below, exploits the joint conditioning of the predicted states of all the agents through the regression feature $\mathbf{z}$ of the GP. Consequently, the control sequence $U_{k}$ explicitly affects the constraint tightening $(\tilde{\mathbb{X}}^{j})$ through Bayesian inference. The dual control effect provides an inherent degree of caution based on the correlation with past observations [3], which is essential to a truly interaction-aware controller. To this end, we make the following proposition.
@@
Algorithm 1 Online Learning GP-MPC
1:  for each time step $k=0,1,2,\dots,$ do
2:     Solve MPC problem (14) for $U^{*}_{k}$
3:     if Problem (14) is feasible, then
4:        Apply $u^{0}_{k}\leftarrow u^{0^{*}}_{0\mid k}$ from (14) to the system
5:     else
6:        Apply shifted input $u^{0}_{k}\leftarrow u_{1\mid k-1}^{0^{*}}$ of the last feasible solution to (14) to the system
7:     end if
8:     Update training set $\mathcal{D}_{k+1}\leftarrow\mathcal{D}_{k}\cup(\mathbf{z}_{k},\mathbf{y}_{k})$
9:     Update inducing points $\mathbf{Z}_{\text{ind}}$ from the solution to (14)
10:  end for
@@ IV-A Lane Merging Use Case
In this section, we demonstrate the proposed GP-MPC method, along with a stochastic baseline MPC, for the motion planning of an autonomous vehicle in a forced lane merging scenario. In this use case, the target lane is occupied by a leading and a following vehicle, as seen in Fig. 1. The Follower tries to close the gap with its Leader and hinders the Ego from merging in between. The Follower’s interactive driving policy is modeled by a Merge-Reactive Intelligent Driver Model (MR-IDM) [16]. We assume the target vehicles remain in their lane center. The Ego, Follower and Leader are denoted by $j=0,1,2,$ respectively. For simplicity and interpretability, we confine ourselves to learning the residual dynamics of a single agent, the Follower $(g^{1})$, assuming the Leader drives at a constant speed $(g^{2}=0)$.
@@ IV-B Vehicle Modeling
All vehicles are modeled by a kinematic bicycle model. The state of a vehicle is described by the state vector $x=\begin{bmatrix}X&Y&v&\psi&\delta\end{bmatrix}^{\top}\in\mathbb{R}^{5},$ where $X$ and $Y$ are the longitudinal and lateral positions of the rear axle, respectively. The longitudinal velocity of the rear axle is denoted by $v$, $\psi$ is the heading angle of the vehicle, and $\delta$ is the steering angle of the front axle. The vehicle’s acceleration $a$ and the steering rate $r$ are inputs to the system: ${u=\begin{bmatrix}a&r\end{bmatrix}^{\top}\in\mathbb{R}^{2}.}$ The continuous-time dynamics of the kinematic bicycle model ${\dot{x}(t)=f_{c}\left(x(t),u\left(t\right)\right)}$ are defined by the state-space equations:
@@
The Follower’s MR-IDM policy [16] is a driver model that considers agents merging into its lane. This closed-loop policy ${a^{1}=a_{\text{MR-IDM}}\bigl{(}\mathbf{x},a^{0},a^{2}\bigr{)}}$ couples the dynamics of the Ego and the Follower. In order to utilize the framework outlined in Section III, we reformulate this policy as a function of $\mathbf{x}_{k}$ and $a^{0}$ through (3). For simulation purposes, we assume the Follower has access to $a^{0}$ and $a^{2}$. The MR-IDM maps the relative longitudinal and lateral position to an effective distance gap for an IDM. The parameters of the Follower’s MR-IDM are listed in Table I. The vehicle inducing the largest deceleration determines the output of the MR-IDM. For details on the MR-IDM we refer the reader to [16].
TABLE I: Merge-Reactive Intelligent Driver Model Parameters.
IDM Parameter Value
Desired speed $(v_{\text{ref}})$ 36 [m/s]
Headway time $(T_{\text{ref}})$ 0.25 [s]
Free acceleration exponent $\left(\delta_{\text{IDM}}\right)$ 4 [-]
Jam distance $\left(s_{0}\right)$ 2 [m]
Maximum acceleration $\left(a_{\text{max}}\right)$ 4 [m/s2]
Desired deceleration $\left(b_{\text{max}}\right)$ 3 [m/s2]
Coolness factor $\left(c\right)$ 0.99 [-]
Lateral reactivity $\left(\zeta\right)$ 1 [-]
Figure 1: Schematic of the lane merging scenario near the merge point.
@@ IV-C Motion Planning Problem
We incentivize the Ego to maintain its initial velocity $v^{0}_{0}$ and stay in its lane, until the merge lane closes and the MPC decides that we should merge. We use a logistic function $m\left(X\right):\mathbb{R}\rightarrow\mathbb{R}$ to describe the center of the merge lane as a function of the longitudinal position, which coincides with the center of the target lane after the merge lane has fully closed. Accordingly, the Ego’s reference $x^{r}$ is defined as:
$x^{r}\left(X\right)=\begin{bmatrix}0&m(X)&v^{0}_{0}&0&0\end{bmatrix}^{\top}.$ (17)
@@
$J$ $\bigl{(}x^{0}_{k},u^{0}_{k-1},U_{k}\bigr{)}=\sum_{i=1}^{N}\left\lVert x^{0}_{i\mid k}-x^{r}\left(X^{0}_{i\mid k}\right)\right\rVert^{2}_{Q}$
$+\sum_{i=0}^{N-1}\left\lVert u^{0}_{i\mid k}\right\rVert^{2}_{R}+\left\lVert\Delta u^{0}_{i\mid k}\right\rVert^{2}_{S},$ (18)
@@
$h_{r}\bigl{(}x^{0}_{i\mid k}\bigr{)}=m(X^{0}_{i\mid k})-Y^{0}_{i\mid k}+\tfrac{W-w}{2}\leq 0,$ (19)
where the road boundary is parallel to the center of the merge lane, $W$ and $w$ denote the vehicle and lane width. Consequently, we have the following general state constraints:
@@
Similar to Zhu et al. [12], we account for the uncertainty in the predicted position of the target vehicle by expanding the semi-axes of elliptical collision avoidance constraints. Again, we assume no uncertainty in their lateral position. By expanding the major semi-axis of the ellipse, we obtain a tractable, tightened reformulation of the chance constraint:
$\begin{split}&\tilde{\mathbb{X}}_{i\mid k}^{j}\bigl{(}x_{i\mid k}^{j}\bigr{)}=\bigl{\{}x^{0}_{i\mid k}\in\mathbb{R}^{n_{x}}\mid h_{c}\bigl{(}x^{0}_{i\mid k},\mu^{x^{j}}_{i\mid k},\Sigma^{X^{j}}_{i\mid k}\bigr{)}=\\ &-\tfrac{\bigl{(}c^{j}_{x,i\mid k}-c^{0}_{x,i\mid k}\bigr{)}^{2}}{\left(\mathcal{E}_{c,A}+\sigma\sqrt{\Sigma^{X^{j}}_{i\mid k}}\right)^{2}}-\tfrac{\bigl{(}c^{j}_{y,i\mid k}-c^{0}_{y,i\mid k}\bigr{)}^{2}}{\mathcal{E}_{c,B}^{2}}+1\leq 0\bigr{\}},\end{split}$ (21)
where $c^{j}_{x,i\mid k}$ and $c^{j}_{y,i\mid k}$ denote the longitudinal and lateral component of the centroid, respectively, and $\Sigma^{X^{j}}_{i\mid k}$ denotes the predicted covariance of the longitudinal position of Agent $j$ at time step $k+i$. The major and minor semi-axis of the ellipse are denoted by $\mathcal{E}_{c,A}$ and $\mathcal{E}_{c,B}$, respectively. The constraint tightening can be tuned with $\sigma\in\mathbb{R}_{\geq 0}$, such that the approximate constraints are enforced with a probability of $p_{x}\geq\text{\normalsize erf}\left(\sigma\right)$, where $\mathrm{erf}$ denotes the Gaussian error function and $\sigma=2$ is the amount of standard deviation that is accounted for. In absence of formal feasibility guarantees of the MPC, the collision avoidance constraints are softened by an $l_{1}$ penalty function to retain feasibility. Since we assume the Leader maintains its initial velocity, we set $\Sigma^{x^{2}}=\boldsymbol{0}$.
@@ IV-D Constant Velocity Model Predictive Control
$v^{j}_{i\mid k}=v^{j}_{k},\quad\text{for }i=0,1,\dots,N.$ (22)
@@
$\mu^{x^{j}}_{i+1\mid k}=A\mu^{x^{j}}_{i\mid k},\quad\text{for }i=0,1,\dots,N-1,\quad j=1,2,$ (23)
@@
The CV-MPC accounts for the Follower’s chance in velocity $\Delta v^{1}$ by a fixed covariance ${\sigma^{2}_{v^{1}}=0.3}$ that we propagate as:
$\Sigma_{i+1\mid k}^{x^{1}}=A\Sigma_{i\mid k}^{x^{1}}+B^{1}\sigma^{2}_{v^{1}},\quad\text{for }i=0,1,\dots,N-1,$ (25)
where $\text{\normalsize where }B^{1}=\begin{bmatrix}0&0&1&0&0\end{bmatrix}^{\top}.$ (26)
@@ IV-E Gaussian Process Model Predictive Control
With the GP-MPC we predict the change velocity using GP regression by measuring the difference between the Follower’s actual velocity and the constant velocity prediction:
@@
The relative lateral position of the Follower and Leader is assumed to be zero and is omitted from $\mathbf{z}$. We use past observations ${\mathbf{y}}$ and approximate the incremental velocity by the posterior of the GP through Bayesian inference:
$d^{1}\bigl{(}\mathbf{z}_{i\mid k}\bigr{)}\mid\mathcal{D},\mathbf{Z}\approx\Delta v^{1}_{i\mid k}=v^{1}_{i+1\mid k}-v^{1}_{i\mid k},$ (29)
and the posterior mean and covariance of the GP affect the predictions via the matrix $B^{1}$ according to (14). Note that the MPC only has access to the current state of other agents.
@@
where $\sigma^{2}_{d}$ is the prior variance and $L_{d}$ denotes the length-scale matrix. If sufficient training data is available, hyper-parameters of the kernel could be inferred using log-likelihood optimization [6]. Here, we use fixed hyper-parameters for the kernel, with $L_{d}=\text{diag}\left(3,3,3,17,17,5\right)$. The covariance of the kernel is equal to the prior covariance of the nominal prediction model $\sigma^{2}_{d}=\sigma^{2}_{\Delta v^{1}}=0.3$, such that the CV-MPC is equivalent to the GP-MPC prior to inference.

### v2

@@ III-C Gaussian Process Prediction Model
We exploit the joint probability of the GP $d^{j}$ to account for the uncertainty in our safety constraint set $\mathbb{X}^{j}$ by formulating chance constraints that should be satisfied with a probability $p_{x}\in\left(0,1\right]$. To this end, we optimize the control sequence of Agent 0 $U_{k}=\bigl{(}u^{0}_{0\mid k},\ldots,u^{0}_{N-1\mid k}\bigr{)}$ in the following SOCP:
@@
Note that the posterior GP $d^{j}(\mathbf{z})$ is a function of the control sequence $U_{k}$. As SOCP (10) is generally intractable, we subsequently apply tractable approximations for consecutive evaluations of the GP and the chance constraints at the expense of formal safety guarantees. In order to propagate the covariance of the predicted state $\Sigma^{x^{j}}$, we make the following assumptions. We assume that consecutive GP evaluations are independent and the predicted state $x^{j}$ and posterior GP $d^{j}(\mathbf{z})$ are assumed to be jointly Gaussian at each prediction step:
$\begin{bmatrix}x^{j}_{i}\\ d^{j}(\mathbf{z}_{i})\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\mu^{x^{j}}_{i}\\ \mu^{d^{j}}(\mathbf{z}_{i})\end{bmatrix},\begin{bmatrix}\Sigma^{x^{j}}_{i}&\Sigma^{x^{j}d^{j}}(\mathbf{z}_{i})\\ \Sigma^{d^{j}x^{j}}(\mathbf{z}_{i})&\Sigma^{d^{j}}(\mathbf{z}_{i})\end{bmatrix}\right),$ (11)
such that the covariance can be approximated and propagated over the horizon. We obtain $\mu^{d^{j}},\Sigma^{x^{j}{d^{j}}},\Sigma^{d^{j}}$ through a first-order Taylor approximation of the GP evaluated at its posterior mean [15], which provides a good trade-off between approximation accuracy and computational complexity [8]:
@@
for $i=0,1,\dots,N-1$, where the initial prediction equals the current state $\mu^{x^{j}}_{0\mid k}=x^{j}_{k}$ with ${\Sigma^{x^{j}}_{0\mid k}=\mathbf{0}}$, and $\Sigma^{j}_{i\mid k}$ denotes the joint covariance matrix (11). We approximate the full GP by a Sparse Pseudo-Input GP (SPGP) [16], which exploits the Nyström projection to reduce the computational complexity of the GP by preconditioning the GP on inducing points. For details on SPGPs, we refer the reader to [16]. We select the inducing points by taking equidistant samples from the predicted trajectory at time $k-1$, since the test points can be expected to be close to this trajectory [8].
@@ III-D Learning-based Dual Model Predictive Control
In this section, we propose our dual MPC that uses the GP prediction model, presented in Sec. III-C, to jointly condition the interactions of the other agents with the control inputs of Agent 00. The dual control effect was first formalized in [3], and according to [4] is defined as follows:
@@
where the chance constraints from (10d) are approximated by a general tightened constraint set $\tilde{\mathbb{X}}^{j}\bigl{(}\mu^{x^{j}},\Sigma^{x^{j}}\bigr{)}\subseteq\mathbb{X}^{j}\bigl{(}\mu^{x^{j}}\bigr{)}$, based on the covariance of the GP [8]. The specific form of constraint tightening may depend on the nature of the problem. An example hereof is provided in Sec. IV. The dual MPC exploits the joint conditioning of the predicted states of all the agents through the regression feature $\mathbf{z}$ of the GP. Consequently, the control sequence $U_{k}$ explicitly affects the constraint tightening $(\tilde{\mathbb{X}}^{j})$ through Bayesian inference. The dual control effect provides an inherent degree of caution based on the correlation with past observations [3], which is essential to a truly interaction-aware controller. To this end, we make the following proposition.
@@ IV-A Lane Merging Use Case
In this section, we demonstrate the proposed GP-MPC method for the motion planning of an autonomous vehicle in a forced lane merging use case, and compare it against a stochastic baseline MPC. This simple use case promotes the qualitative interpretability of GP-MPC. Here, the target lane is occupied by a leading and a following vehicle, as seen in Fig. 1 in Sec. V. The Ego vehicle is controlled by the GP-MPC from (14). The Follower’s interactive driving policy follows a Merge-Reactive Intelligent Driver Model (MR-IDM) [17]. The Follower tries to close the gap with its Leader, hindering the Ego from merging. We assume the target vehicles remain in their lane center. The Ego, Follower, and Leader are denoted by $j=0,1,2$, respectively. For simplicity, we learn only the residual dynamics of the Follower $(g^{1})$, assuming the Leader drives at constant speed $(g^{2}=0)$, and assume the disturbance processes $w^{j}_{k}$ are zero.
@@ IV-B Vehicle Modeling
All vehicles are modeled by a deterministic, kinematic bicycle model, which is sufficiently accurate for motion planning. The state of a vehicle is described by the state vector $x=\begin{bmatrix}X&Y&v&\psi&\delta\end{bmatrix}^{\top}\in\mathbb{R}^{5},$ where $X$ and $Y$ are the longitudinal and lateral positions of the rear axle, respectively. The longitudinal velocity of the rear axle is denoted by $v$, $\psi$ is the heading angle of the vehicle, and $\delta$ is the steering angle of the front axle. The vehicle’s acceleration $a$ and the steering rate $r$ are inputs to the system: ${u=\begin{bmatrix}a&r\end{bmatrix}^{\top}\in\mathbb{R}^{2}}$. The continuous-time dynamics of the kinematic bicycle model ${\dot{x}(t)=f_{c}\left(x(t),u\left(t\right)\right)}$ are defined by the state-space equations:
@@
The Follower’s MR-IDM policy [17] is a driver model that considers agents merging into its lane. This closed-loop policy ${a^{1}=a_{\text{MR-IDM}}\bigl{(}\mathbf{x},a^{0},a^{2}\bigr{)}}$ couples the dynamics of the Ego and the Follower. In order to utilize the framework outlined in Sec. III, we reformulate this policy as a function of $\mathbf{x}_{k}$ and $u^{0}$ through (3). For simulation purposes, we assume the Follower has access to $a^{0}$ and $a^{2}=0$. The MR-IDM maps the relative longitudinal and lateral position to an effective distance gap for an IDM. The vehicle inducing the largest deceleration determines the output of the MR-IDM. For details on the MR-IDM we refer the reader to [17].
@@ IV-C Motion Planning Problem
We incentivize the Ego to maintain its initial velocity $v^{0}_{0}$ and stay in its lane, until the merge lane closes and the MPC decides that we should merge. We use a smooth function $m\left(X\right):\mathbb{R}\rightarrow\mathbb{R}$ to describe the center of the merge lane as a function of the longitudinal position, which coincides with the center of the target lane after the merge lane has fully closed. Accordingly, the Ego’s reference $x^{r}$ is defined as:
$x^{r}\left(X\right)=\begin{bmatrix}0&m(X)&v^{0}_{0}&0&0\end{bmatrix}^{\top}.\vspace{-1mm}$ (17)
@@
$J$ $\bigl{(}x^{0}_{k},u^{0}_{k-1},U_{k}\bigr{)}=\sum_{i=1}^{N}\left\lVert x^{0}_{i\mid k}-x^{r}\left(X^{0}_{i\mid k}\right)\right\rVert^{2}_{Q}$ (18)
$+\sum_{i=0}^{N-1}\left\lVert u^{0}_{i\mid k}\right\rVert^{2}_{R}+\left\lVert\Delta u^{0}_{i\mid k}\right\rVert^{2}_{S},$
@@
$h_{r}\bigl{(}x^{0}_{i\mid k}\bigr{)}=m\bigl{(}X^{0}_{i\mid k}\bigr{)}-Y^{0}_{i\mid k}+\tfrac{W-w}{2}\leq 0,\vspace{-1mm}$ (19)
where the road boundary is parallel to the center of the merge lane, $W$ and $w$ denote the vehicle and lane width. Consequently, we have the following state constraint set:
@@
Similar to Zhu et al. [13], we account for the uncertainty in the predicted position of the target vehicle by expanding the semi-axes of elliptical collision avoidance constraints. Again, we assume no uncertainty in their lateral position. By expanding the major semi-axis of the ellipse, we obtain a tractable, tightened reformulation of the chance constraints:
$\tilde{\mathbb{X}}_{i\mid k}^{j}\bigl{(}x_{i\mid k}^{j}\bigr{)}=\bigl{\{}x^{0}_{i\mid k}\in\mathbb{R}^{n_{x}}\mid h_{c}\bigl{(}x^{0}_{i\mid k},\mu^{x^{j}}_{i\mid k},\Sigma^{X^{j}}_{i\mid k}\bigr{)}=$ (21)
$-\tfrac{\bigl{(}c^{j}_{x,i\mid k}-c^{0}_{x,i\mid k}\bigr{)}^{2}}{\left(\mathcal{E}_{c,A}+\sigma\sqrt{\Sigma^{X^{j}}_{i\mid k}}\right)^{2}}-\tfrac{\bigl{(}c^{j}_{y,i\mid k}-c^{0}_{y,i\mid k}\bigr{)}^{2}}{\mathcal{E}_{c,B}^{2}}+1\leq 0\bigr{\}},$
where $c^{j}_{x,i\mid k}$ and $c^{j}_{y,i\mid k}$ denote the longitudinal and lateral component of the centroid, respectively, and $\Sigma^{X^{j}}_{i\mid k}$ denotes the predicted covariance of the longitudinal position of Agent $j$ at time step $k+i$. The major and minor semi-axis of the ellipse are denoted by $\mathcal{E}_{c,A}$ and $\mathcal{E}_{c,B}$, respectively. The constraint tightening can be tuned with $\sigma\in\mathbb{R}_{\geq 0}$, where $\sigma=2$ is the amount of standard deviation that is accounted for. In absence of formal feasibility guarantees of the MPC, the collision avoidance constraints are softened using an $l_{1}$ penalty function to retain feasibility. Since we assume the Leader maintains its initial velocity, we set $\Sigma^{x^{2}}=\boldsymbol{0}$.
@@ IV-D Constant Velocity Model Predictive Control
$v^{j}_{i\mid k}=v^{j}_{k},\quad\text{for }i=0,1,\dots,N.\vspace{-1mm}$ (22)
@@
$\mu^{x^{j}}_{i+1\mid k}=A\mu^{x^{j}}_{i\mid k},\quad\text{for }i=0,1,\dots,N-1,\quad j=1,2,\vspace{-2mm}$ (23)
@@
The CV-MPC accounts for the Follower’s change in velocity $\Delta v^{1}$ by a fixed covariance ${\sigma^{2}_{v^{1}}=0.3}$ that we propagate as:
$\Sigma_{i+1\mid k}^{x^{1}}=A\Sigma_{i\mid k}^{x^{1}}A^{\top}+B^{1}\sigma^{2}_{v^{1}}B^{1\top},\quad$ (25)
for $i=0,1,\dots,N-1$, where $B^{1}=\begin{bmatrix}0&0&1&0&0\end{bmatrix}^{\top}$.
@@ IV-E Gaussian Process Model Predictive Control
We use past observations ${\mathbf{y}}$ and predict the incremental velocity by the posterior of the GP through Bayesian inference. Recall that the MPC only has access to the current state of other agents. Therefore, the GP-MPC predicts the change of velocity by observing the difference between the Follower’s actual velocity and the constant velocity prediction:
@@
The relative lateral position of the Follower and Leader is assumed to be zero and is omitted from training point $\mathbf{z}$.
@@
where $\sigma^{2}_{d}$ is the prior covariance and $L_{d}$ is the length-scale matrix. If sufficient training data is available, these hyper-parameters can be inferred using log-likelihood optimization [6]. Here, we use fixed hyper-parameters for the kernel with $L_{d}=\text{diag}\left(3,3,3,17,17,5\right)$. The prior covariance is equal to the covariance of the baseline prediction model $\sigma^{2}_{d}=0.3$, making the CV-MPC equivalent to the GP-MPC prior to inference. Since we do not consider any noise, we set $\sigma^{2}_{w}=0$.